# Quasi-randomization tests for network interference

### v1

@@ Abstract
Many classical inferential approaches fail to hold when interference exists among the population units. This amounts to the treatment status of one unit affecting the potential outcome of other units in the population. Testing for such spillover effects in this setting makes the null hypothesis non-sharp. An interesting approach to tackling the non-sharp nature of the null hypothesis in this setup is constructing conditional randomization tests such that the null is sharp on the restricted population. In randomized experiments, conditional randomized tests hold finite sample validity. Such approaches can pose computational challenges as finding these appropriate sub-populations based on experimental design can involve solving an NP-hard problem. In this paper, we view the network amongst the population as a random variable instead of being fixed. We propose a new approach that builds a conditional quasi-randomization test. Our main idea is to build the (non-sharp) null distribution of no spillover effects using random graph null models. We show that our method is exactly valid in finite-samples under mild assumptions. Our method displays enhanced power over other methods, with substantial improvement in complex experimental designs. We highlight that the method reduces to a simple permutation test, making it easy to implement in practice. We conduct a simulation study to verify the finite-sample validity of our approach and illustrate our methodology to test for interference in a weather insurance adoption experiment run in rural China.
Keywords: Interference, Networks, Causal Inference, Quasi-Randomization Tests, Peer Effects, Random Graphs
@@ 1   Introduction
A common assumption in causal inference is that of the Stable Unit Treatment Value Assumption (SUTVA). This assumes away any dependence of one unit’s treatment status on the outcome of another unit. Whilst this allows for unbiased treatment effect estimation and valid inferential strategies, there are many experimental and observational settings where interference between units takes place (Cox (1958)). There are multiple pathways via which interference can take place in the population. It can be a result of the spread of treatment amongst the population units, the respective outcomes of treated units affecting the outcomes of others, or a combination of both. For example, the spread of a contagious disease is contingent upon the number of infected people in the population (Halloran and Struchiner (1995)). Thus, the estimation of any preventive measures taken will be challenged because of a likely SUTVA violation. Consider a population-level policy evaluation program, like an insurance product campaign (Cai et al. (2015)). Due to social network ties in the population, information dissemination is a plausible interference mechanism that can occur and affect evaluated estimates. Similar phenomena can be seen amongst experiments run on social media networks, like Linkedin, or online marketplaces like eBay, owing to the interconnected nature of these platforms (Pouget-Abadie et al. (2019); Blake and Coey (2014)). The total treatment effect, in such cases, can be decomposed into two effects, one of the direct effect and the other of spillover effects (Sobel (2006)).
@@
The classical method for testing treatment effects is the “Fisher Randomization Test” or FRT (Fisher (1936)). It is used to test for a sharp null hypothesis of no treatment effect for all units in the population. The test exploits the imputability of counterfactuals under the null. This enables the determination of the randomization null distribution of a test statistic of choice. If one were to test for spillover effects, it would render the null non-sharp, leading to imputability issues for the test statistic. There is a growing body of literature that addresses this issue by considering a carefully determined subset of units in the population and treatment assignment vector, such that on these selected units and treatment vectors, the null is sharp (Aronow and Samii (2017); Athey et al. (2018); (Basse et al., 2019b)). Such conditional randomization tests also lead to valid exact p-values, as is the case with the Fisher Randomization Test. Basse et al. (2019b) build a permutation test under the special case of clustered interference. Whilst a permutation test is computationally cheaper, the method lacks generalizability because of the structured nature of interference considered. Aronow and Samii (2017) and Athey et al. (2018) generalize the above for a given network and consider a range of non-sharp null hypotheses for testing. Puelz et al. (2022) presents a graph-theoretic approach to constructing the conditioning mechanism of a sub-population of units and the corresponding treatment assignment vectors for which the null is sharp. However, the proposed methodology is NP-hard in implementation and, thus, in practice, does not involve observing the treatment assignment to construct the conditioning mechanism. Whilst the above considers the problem of testing for interference in an experimental setting, the literature on testing for spillover effects in an observational setting is still in its infancy.
@@
In this paper, we consider the network to be coming from a distribution itself. This is in contrast with the setup considered in the current methods that assume the network to be fixed. We draw a parallel here to estimating the probability of unit treatment assignment via propensity scores in the case of observational studies. The propensity of two nodes in a network being connected can be looked at as being a function of their network characteristics, covariates, or both. This leads us to think of these linkages as coming from a generating process with an associated probability distribution. The current methods are based on strict randomization inference with no distributional assumptions for either the network or the potential outcomes, achieving a generalization of the Fisher Randomization Test. Whilst this allows for widespread validity of the methods proposed, it comes at a cost of loss of power. We emphasize that leveraging appropriate stochastic assumptions can significantly boost power and lead to methodological advancements. In this paper, we restrict the network data-generating process to be characterized by any arbitrary degree distribution. First proposed in Newman et al. (2001), random graphs with arbitrary degree distribution represent real-world social networks closely owing to their ability to capture heterogeneity in degree. Given the mild nature of the distributional assumption, we obtain a large class of network data-generating processes. This assists us in building a quasi-randomization test based on the network distribution. We show that conditional on the observed treatment, the (quasi) randomization distribution of a test statistic leads to valid p-values. The challenge of the imputability of a test statistic of choice under the null, being the major methodological contribution of the current methods in the literature, is naturally overcome in our method.
@@
Model-based approaches with distribution assumptions on outcome model function have been previously considered in the literature for identifying and estimating peer effects (Manski (1993); Manski (2013); Bowers et al. (2013); Toulis and Kao (2013); Blume et al. (2015)). More recently, Li and Wager (2022) and Leung (2020) proposed estimation strategies for spillover effects under a distributional assumption on the network formation model. The inferential methods for the proposed estimation procedures involve either non-standard procedures or asymptotic approximations (Imbens and Manski (2004)). With a careful identification of causal estimates, such large sample normal approximations of confidence interval conform to a Neymanian perspective. Our first contribution is to extend the random graph interference model into the potential outcome framework. We generalize and define sharp/non-sharp null hypotheses for random graph models. Building upon this, we develop a randomization-based test for a given random graph model. This leads to finite sample validity of the p-value obtained for the oracle method, in line with Fisherian procedures.
@@
Our method addresses several technical challenges encountered in the current methods. We show that randomization tests based on network-generating processes lead to imputable test statistics under the null, conditional on the treatment assignment. This overcomes a major challenge faced in randomization methods based on random treatment assignment. The current methodology involves restricting attention to a carefully chosen subset of focal units and claims null of no spillover effects for only these units. In contrast, our method tests for a stricter class of nulls with no spillover effects on all the units in line with the Fisher Randomization Test. Another hurdle faced by the current methodologies is the extension of the method to complex experimental designs. Determination of conditional randomization distribution given a conditioning mechanism is not straightforward. The restriction of fixed focal units in the randomization null reduces the sample space of possible treatment assignment vectors in the null. It can lead to an enormous loss of power or render the null degenerate completely. There is a burgeoning set of literature focussing on improving the current methods by incorporating design-based techniques (Puelz et al. (2022); (Basse et al., 2019a)). We emphasize that our method is easily generalized to various experimental designs, and we show this for two experimental designs, namely, completely randomized experiments and cluster randomized experiments.
@@
Our second major contribution is the computational efficiency of the proposed method. We propose that the random graph model be characterized by any arbitrary degree distribution. We build a permutation test by generating the null over the set of graphs with the same degree sequence as the observed graph. Random graph generation of a prescribed degree sequence admits fast and efficient polynomial time algorithms for large graphs (Viger and Latapy (2005); Fosdick et al. (2018)). Thus, via permutation of edges of the graph in a manner that preserves degree sequence, we obtain valid randomization tests. This is closely related to literature in random graph null model used in network analysis (Van Koevering et al. (2021)) with applications in protein structure discovery and community detection among many (Milenković et al. (2009); Sah et al. (2014)). To the best of our knowledge, we are the first to formally introduce random graph null models as an extension of the Fisher Randomization Test. To illustrate our method, we re-analyze a completely randomized experiment in rural China where rice crop farmers are informed about a weather insurance product (Cai et al. (2015)). The hypothesis is that farmers with a higher proportion of informed neighbors display an enhanced adoption rate. We display our methodology on a subset of the original experiment performed and obtain significant information spillover among the farmers. The setting displays the importance of such spillover effects in evidence-based policymaking under resource-constrained settings.
@@
The remaining paper is structured as follows. In Section 2, we set up notation and formalize the null hypothesis of interest, incorporating network random variable. We also describe the present state-of-the-art methodology of using a conditioning mechanism for testing non-sharp null. Section 3 presents a randomization test based on a given network-generating process and its application to different experimental designs. Section 4 extends this further to random graph models and presents our main methodology, which can be implemented as a permutation test. In Section 5, we present extensive numerical simulations to evaluate and compare the power of the proposed test with the current methods in the literature. Section 6 and 7 illustrate our methodology as an empirical application and conclude the paper. All theorem proofs, algorithms, and additional numerical simulation results are presented in the Appendix.
@@ 2.1   Notation
Consider a population =$\mathbb{P}=\{1,2,...,N\}$ where $N$ is the total number of units in the population. Consider a vector $Z\in{\mathbb{Z}}^{N}$ where each element in the vector indicates treatment assigned to the corresponding indexed unit in the population. Here, $\mathbb{Z}$ represents the set of possible treatments to which the population units can be exposed. For the purposes of this article, we take =$\mathbb{Z}=\{0,1\}$. Hence, any treatment assignment vector is an N-dimensional binary vector. Consider the treatment vector to be drawn from a probability distribution $P_{Z}:\{0,1\}^{N}\rightarrow[0,1]\subset\mathbb{R}$. Let $Z_{obs}$ be the realized outcome drawn from the treatment assignment distribution $P_{Z}$.
@@
We consider the population $\mathbb{P}$ to be connected via an undirected graph $G=(\mathbb{P},E)$ where $E\subseteq\mathbb{P}\times\mathbb{P}$. For distinct units $i$ and $j$ such that $(i,j)\in E$ represents a connection between the two units. With some slight abuse of notation, we also represent the adjacency matrix of the graph by $G$. Thus, $G$ is an $N\times N$ binary matrix. Here, for some $i$ and $j$, $G_{ii}=0$ and $G_{ij}=G_{ji}$ owing to $G$ being a undirected graph with no self-loops. If $G_{ij}=1$, then we say unit $i$ and $j$ are neighbours. This can also be denoted by $i\in\mathcal{N}_{G}(j)$ or vice-versa. Here, $\mathcal{N}_{G}(i)$ represents all the neighbors of unit $i$ in graph $G$. We may simply write $\mathcal{N}(i)$ if the graph is clear in context. We define the total number of neighbors of unit $i$ as the degree of $i$. This is denoted by $deg(i)$ and is equal to $|N(i)|$. We denote the distance between two units $i$ and $j$ as $dist(i,j)$. It is defined as the shortest path length between the two units. By convention, we take $dist(i,i)=0$ and $dist(i,j)=\infty$ if no path exists between the two distinct units. Consider the graph $G$ to be coming from the probability distribution $P_{G}:\{0,1\}^{N\times N}\rightarrow[0,1]\subset\mathbb{R}$. We denote the observed sample from the distribution $P_{G}$ to be $G_{obs}$. Each unit $i$ in the population is ‘characterized’ by a set of pre-treatment or covariate vectors given by $X_{i}$ where $X_{i}$ is a m-dimensional vector. $X_{i}$ is fixed and influences the outcome via the potential outcome function. Let $X$ denote the $N\times M$ matrix where each row corresponds to the respective unit’s covariate vector. We define the potential outcome function $Y:\{0,1\}^{N}\times\{0,1\}^{N\times N}\rightarrow\mathbb{R}^{N}$ as a real-valued function that takes treatment assignment and network adjacency matrix as arguments. We will consider $Y$ to be fixed, and each element $i$ of an image in $Y(Z,G)$, $Y_{i}(Z,G)$ denotes the outcome of $i^{th}$ unit in the population under the treatment assignment $Z$ and network assignment $G$. We denote the corresponding potential outcome as $Y_{obs}$ that is, $Y_{obs}=Y(Z_{obs},G_{obs})$. We note that the potential outcome function implicitly contains information in $X$ and can be written as $Y_{X}(Z,G)$. For the remainder of the article, we will use the notation $Y(Z,G)$ for readability.
@@ 2.2   Definitions
The assignment mechanism defines the joint probability distribution of treatment assignment vectors given covariates and potential outcomes, denoted by $P(Z|X,Y(0),Y(1))$. The following are three key assumptions that are considered in the causal inference literature for assignment mechanism: (i) individualistic assignment, which restricts the treatment assignment of a unit to its own covariates and outcomes (ii) probabilistic assignment, which assumes non-integer probabilities of treatment assignment to units (iii) unconfoundedness assumption, which restricts the dependence of the probability assignment of units to only covariates, and not the potential outcome (Imbens and Rubin (2015)). Individualistic assignment mechanism assumption, given a well-defined potential outcome function, constitutes Stable Unit Treatment Value Assumption (SUTVA). We argue that the underlying network structure among the population is an implicit form of treatment. Hence, we generalize the treatment to a bivariate treatment assignment vector $(Z,G)$. Whilst the usual assumptions described above are plausible for the marginal distribution of direct treatment $Z$, the individualistic assumption is not valid for the marginal distribution of network treatment $G$ given the nature of the variable. We look at the distribution of joint treatment assignment vector $(Z,G)$ in this setup.
$P(Z,G|X)=P(G|Z,X)\cdot P(Z|X),$
assuming joint treatment assignment is unconfounded, that is,
@@
We remark that we took the assumption of the unconfoundedness of both treatment assignments into consideration. We formally define the standard sharp null hypothesis of no (direct) treatment effect when the population is not interconnected.
@@
Sharp null hypothesis of no treatment effect:
@@
The above stated null hypothesis imposes specified forms of restrictions on the potential outcome function. Extending this to our setup, we get the following definition of the sharp null hypothesis.
@@
Sharp null hypothesis of no (joint) treatment effect:
@@
We now give a definition of a test statistic that will be used for testing such null hypotheses. Following our setup, we will use the joint treatment assignment, $(Z,G)$, for the rest of the article.
@@
Under sharp null, given an observed outcome, $Y_{obs}$, we can impute the outcome function $Y$ for all treatment assignments. This is not possible when the null hypothesis is non-sharp. We formally define the non-sharp null hypothesis below.
@@
Here, the subset restriction, denoted by $\textbf{Z}\times\textbf{G}:=\textbf{Z}_{i}\times\textbf{G}_{i}$, on treatment assignment is chosen in accordance with the null hypothesis of interest. We elaborate further using the example of testing null of no spillover effects below. Consider the null hypothesis of no spillover effects if the distance between two units is at least k.
@@
$\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})\qquad$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
[…]
Note the non-imputability of potential outcome function under the null considered above for different treatment assignments. Specifically, given $(Z_{obs},G_{obs})$, we can compute potential outcomes for treatment assignments in the restriction set defined above. Since the restriction set is a strict subset of the treatment assignment space, the null is not sharp.

### v2

@@ Abstract
Network interference amounts to the treatment status of one unit affecting the potential outcome of other units in the population. Testing for spillover effects in this setting makes the null hypothesis non-sharp. An interesting approach to tackling the non-sharp nature of the null hypothesis in this setup is constructing conditional randomization tests such that the null is sharp on the restricted population. Such approaches can pose computational challenges as finding these appropriate sub-populations based on experimental design can involve solving an NP-hard problem. In this paper, we view the network amongst the population as a random variable instead of being fixed. We propose a new approach that builds a conditional quasi-randomization test. We build the (non-sharp) null distribution of no spillover effects using random graph null models. We show that our method is exactly valid in finite samples under mild assumptions. Our method displays enhanced power over other methods, substantially improving cluster randomized trials. We illustrate our methodology to test for interference in a weather insurance adoption experiment run in rural China.
Keywords: Causal Inference, Network Interference, Peer Effects, Randomization Tests, Random Graphs
@@ 1 Introduction
A common assumption in causal inference is that of the Stable Unit Treatment Value Assumption (SUTVA). SUTVA assumes no dependence of a unit’s outcome on another unit’s treatment status. While this allows for unbiased treatment effect estimation and valid inferential strategies, there are many experimental and observational settings where interference between units takes place (Cox (1958)). There are multiple pathways via which interference can take place in the population. It can result from the spread of treatment amongst the population units, the respective outcomes of treated units affecting the outcomes of others, or a combination of both. For example, the spread of a contagious disease depends on the number of infected people in the population (Halloran and Struchiner (1995)). Thus, the estimation of any preventive measures taken will be challenged because of a likely SUTVA violation. Consider a population-level policy evaluation program, like an insurance product campaign (Cai et al. (2015)). Due to social network ties in the population, information dissemination is a plausible interference mechanism that can occur and affect evaluated estimates. Similar phenomena can be seen amongst experiments run on social media networks, like Linkedin, or online marketplaces like eBay, owing to the interconnected nature of these platforms (Pouget-Abadie et al. (2019); Blake and Coey (2014)). The total treatment effect, in such cases, can be decomposed into two effects, one of the direct effects and the other of spillover effects (Sobel (2006)).
@@
The classical method for testing treatment effects is the “Fisher Randomization Test” or FRT (Fisher (1936)). FRT is used to test for a sharp null hypothesis of no treatment effect for all units in the population. The test exploits the imputability of counterfactuals under the null. Imputation enables the determination of the randomization null distribution of a test statistic of choice. If one were to test for spillover effects, it would render the null non-sharp, leading to imputability issues for the test statistic. There is a growing body of literature that addresses this issue by considering a carefully determined subset of units in the population and treatment assignment vector, such that on these selected units and treatment vectors, the null is sharp (Aronow and Samii (2017); Athey et al. (2018); Basse et al. (2019b)). Such conditional randomization tests also lead to valid exact p-values, as with the Fisher Randomization Test. Basse et al. (2019b) build a permutation test under the particular case of clustered interference. While a permutation test is computationally cheaper, the method lacks generalizability because of the structured nature of the interference considered. Aronow and Samii (2017) and Athey et al. (2018) generalize the above for a given network and evaluate a range of non-sharp null hypotheses for testing. Puelz et al. (2022) presents a graph-theoretic approach to constructing the conditioning mechanism of a sub-population of units and the corresponding treatment assignment vectors for which the null is sharp. However, the proposed methodologies lose power in cluster randomized experiments.
@@
This paper considers the network to be coming from a distribution itself. This idea contrasts with the setup considered in the current methods, assuming the network is fixed. We draw a parallel here to estimating the probability of unit treatment assignment via propensity scores in the case of observational studies. The propensity of two nodes in a network being connected can be considered a function of their network characteristics, covariates, or both. This leads us to think of these linkages as coming from a generating process with an associated probability distribution. The current methods are based on strict randomization inference with no distributional assumptions for either the network or the potential outcomes, achieving a generalization of the Fisher Randomization Test. This allows for the widespread validity of the methods proposed, albeit it might be at the cost of power. This paper explores how leveraging appropriate stochastic assumptions can significantly boost power and lead to methodological advancements.
We assume that a given vector of sufficient statistics characterizes the network data-generating process and can thus be represented as an exponential random graph model. We first take the degree sequence of the graph as a sufficient statistic. Proposed in Newman et al. (2001), random graphs with arbitrary degree distribution represent real-world social networks closely owing to their ability to capture heterogeneity in degree. This assists us in building a quasi-randomization test based on the network distribution by randomizing over graphs from this generating process. We show that conditional on the observed treatment, a test statistic’s (quasi) randomization distribution leads to valid p-values. We then generalize the model to a vector of sufficient statistics and present a method that works for any set of such network statistics.
@@
Model-based approaches with distribution assumptions on outcome model function have been previously considered in the literature for identifying and estimating peer effects (Manski (1993, 2013); Bowers et al. (2013); Toulis and Kao (2013); Blume et al. (2015)). More recently, Li and Wager (2022) and Leung (2020) proposed estimation strategies for spillover effects under a distributional assumption on the network formation model. The inferential methods for the proposed estimation procedures involve non-standard or asymptotic approximations (Imbens and Manski (2004)). With a careful identification of causal estimates, such large sample normal approximations of confidence interval conform to a Neymanian perspective.
Our first contribution is to extend the random graph interference model into the potential outcome framework. We generalize and define sharp/non-sharp null hypotheses for random graph models. Building upon this, we develop a randomization-based test for a given random graph model. This leads to finite sample validity of the p-value obtained, in line with Fisherian procedures.
@@
As a second contribution, our method addresses several technical challenges encountered in the current methods. We show that randomization tests based on network-generating processes lead to imputable test statistics under the null, conditional on the treatment assignment. This overcomes a significant challenge in randomization methods based on random treatment assignment. The state-of-the-art techniques restrict attention to a carefully chosen subset of focal units and claim null of no spillover effects for only these units. In contrast, our method tests for a stricter class of nulls with no spillover effects on all the units.
Another hurdle the current methodologies face is the loss of power in cluster randomized trials. Determination of conditional randomization distribution given a conditioning mechanism is not straightforward. The restriction of fixed focal units in the randomization null reduces the sample space of possible treatment assignment vectors in the null. It can lead to an enormous loss of power or render the null degenerate completely. A burgeoning literature set focuses on improving current methods by incorporating design-based techniques (Puelz et al. (2022); Basse et al. (2019a)). Unlike distribution-free methods, we show how incorporating information about network generation into the methodology can boost power.
@@
Our third significant contribution is the computational efficiency of the proposed method. We suggest that sufficient network statistics can characterize the random graph model. We first take a single sufficient statistic as the network’s degree sequence and build a permutation test by generating the null over the set of graphs with the same degree sequence as the observed graph. Random graph generation of a prescribed degree sequence admits fast and efficient polynomial time algorithms for large graphs (Viger and Latapy (2005); Fosdick et al. (2018)). Thus, we obtain valid randomization tests via permutation of edges of the graph in a manner that preserves degree sequence. We then present a method for a general exponential random graph model characterized by an arbitrary set of network statistics and build a permutation test by generating graphs in the same isomorphism class under the null. This is closely related to the literature on the random graph null model used in network analysis with applications in protein structure discovery and community detection among many (Milenković et al. (2009); Sah et al. (2014)).
To illustrate our method, we re-analyze a randomized experiment in rural China where rice crop farmers are informed about a weather insurance product (Cai et al. (2015)). The hypothesis is that farmers with a higher proportion of informed neighbors display an enhanced adoption rate. We display our methodology on a subset of the original experiment performed and obtain significant information spillover among the farmers.
@@
The remaining paper is structured as follows. In Section 2, we set up notation and formalize the null hypothesis of interest, incorporating network random variable. Section 3 presents a randomization test based on a given network-generating process and its application to different experimental designs. Section 4 extends this further to random graph models and presents our central methodology introducing the permutation test. In Section 5, we present extensive numerical simulations to evaluate the performance of the proposed test. Sections 6 and 7 illustrate our methodology as an empirical application and conclude the paper. Proofs, algorithms, and additional numerical results are presented in the Appendix.
@@ 2.1 Notation
Consider a population =$\mathbb{P}=\{1,2,...,N\}$ where $N$ is the total number of units in the population. Consider a vector $Z\in{\mathbb{Z}}^{N}$ where each element in the vector indicates treatment assigned to the corresponding indexed unit in the population. Here, $\mathbb{Z}$ represents the set of possible treatments to which the population units can be exposed. We take =$\mathbb{Z}=\{0,1\}$ for this article. Hence, any treatment assignment vector is an N-dimensional binary vector. We consider the population $\mathbb{P}$ to be connected via an undirected graph $G=(\mathbb{P},E)$ where $E\subseteq\mathbb{P}\times\mathbb{P}$. For distinct units $i$ and $j$ such that $(i,j)\in E$ represents a connection between the two units. With some slight abuse of notation, we also represent the graph’s adjacency matrix by $G$. Thus, $G$ is an $N\times N$ binary matrix. Here, for some $i$ and $j$, $G_{ii}=0$ and $G_{ij}=G_{ji}$ owing to $G$ being a undirected graph with no self-loops. If $G_{ij}=1$, then we say unit $i$ and $j$ are neighbours. This can also be denoted by $i\in\mathcal{N}_{G}(j)$ or vice-versa. Here, $\mathcal{N}_{G}(i)$ represents all the neighbors of unit $i$ in graph $G$. We may write $\mathcal{N}(i)$ if the graph is clear in context. We define the total number of unit neighbors $i$ as the degree of $i$. This is denoted by $deg(i)$ and equals $|N(i)|$. We denote the distance between two units $i$ and $j$ as $dist(i,j)$. It is defined as the shortest path length between the two units. By convention, we take $dist(i,i)=0$ and $dist(i,j)=\infty$ if no path exists between the two distinct units. Consider $\mathbb{P^{\prime}}\subseteq\mathbb{P}$. We denote $G[\mathbb{P^{\prime}}]$ as the sub-graph induced by $\mathbb{P^{\prime}}$. That is, $G[\mathbb{P^{\prime}}]=(\mathbb{P^{\prime}},E^{\prime})$ where $E^{\prime}=\{(i,j)\in\mathbb{P^{\prime}}\times\mathbb{P^{\prime}}|\>(i,j)\in E\}$.
Each unit $i$ in the population is ‘characterized’ by a set of pre-treatment or covariate vectors given by $X_{i}$ where $X_{i}$ is a m-dimensional vector. $X_{i}$ is fixed and influences the outcome via the network’s potential outcome function or formation. Let $X$ denote the $N\times M$ matrix where each row corresponds to the respective unit’s covariate vector. Consider the treatment vector to be drawn from a probability distribution $P_{X}(Z,G):\{0,1\}^{N}\times\{0,1\}^{N\times N}\rightarrow[0,1]$. We define the conditional probability distribution $P(Z|G):\{0,1\}^{N}\rightarrow[0,1]\subset\mathbb{R}$ as $P_{Z|G}$. Let $Z_{obs}$ be the realized treatment drawn from the treatment assignment distribution $P_{Z|G}$. Consider the graph $G$ to be coming from the marginal probability distribution $P(G):\{0,1\}^{N\times N}\rightarrow[0,1]\subset\mathbb{R}$, denoted as $P_{G}$. We denote the observed sample from the distribution $P_{G}$ to be $G_{obs}$. We define the potential outcome function $Y:\{0,1\}^{N}\times\{0,1\}^{N\times N}\rightarrow\mathbb{R}^{N}$ as a real-valued function that takes treatment assignment and network adjacency matrix as arguments. We will consider $Y$ to be fixed, and each element $i$ of an image in $Y(Z,G)$, $Y_{i}(Z,G)$ denotes the outcome of $i^{th}$ unit in the population under the treatment assignment $Z$ and network assignment $G$. We denote the corresponding potential outcome as $Y_{obs}$ that is, $Y_{obs}=Y(Z_{obs},G_{obs})$. We note that the potential outcome function implicitly contains information in $X$ and can be written as $Y_{X}(Z,G)$. For the remainder of the article, we will use the notation $Y(Z,G)$ for readability.
@@ 2.2 Definitions
The assignment mechanism defines the joint probability distribution of treatment assignment vectors given covariates and potential outcomes, denoted by $P(Z|X,Y(0),Y(1))$. The following are three key assumptions that are considered in the causal inference literature for assignment mechanism: (i) individualistic assignment, which restricts the treatment assignment of a unit to its covariates and outcomes (ii) probabilistic assignment, which assumes non-integer probabilities of treatment assignment to units (iii) unconfoundedness assumption, which restricts the dependence of the probability assignment of units to only covariates, and not the potential outcome (Imbens and Rubin (2015)). Given a well-defined potential outcome function, individualistic assignment mechanism assumption constitutes Stable Unit Treatment Value Assumption (SUTVA). We argue that the underlying network structure among the population is an implicit form of treatment. Hence, we generalize the treatment to a bivariate treatment assignment vector $(Z,G)$. We assume that the joint treatment assignment is unconfounded; that is,
@@
We formally define the standard sharp null hypothesis of no (direct) treatment effect when the population is not interconnected.
@@
The sharp null hypothesis of no treatment effect:
@@
The above-stated null hypothesis imposes specified restrictions on the potential outcome function. Extending this to our setup, we get the following definition of the sharp null hypothesis.
@@
The sharp null hypothesis of no (joint) treatment effect:
@@
We now define a test statistic that will be used for testing such null hypotheses. Following our setup, we will use the joint treatment assignment, $(Z,G)$, for the rest of the article.
@@
Under the sharp null, given an observed outcome, $Y_{obs}$, we can impute the outcome function $Y$ for all treatment assignments. This is not possible when the null hypothesis is non-sharp. We formally define the non-sharp null hypothesis below.
@@
Here, the subset restriction, denoted by $\textbf{Z}\times\textbf{G}:=\textbf{Z}_{i}\times\textbf{G}_{i}$, on treatment assignment is chosen by the null hypothesis of interest. We elaborate further using the testing null of no spillover effects example below. Consider the null hypothesis of no spillover effects if the distance between two units is at least k.
@@
$D_{G}^{k}(i)=\begin{cases}1&\text{if dist(i,j)}\leq k,\\ 0&\text{otherwise}.\end{cases}$
$\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
@@
Note the non-imputability of the potential outcome function under the null considered above for different treatment assignments. Precisely, given $(Z_{obs},G_{obs})$, we can compute potential outcomes for treatment assignments in the restriction set defined above. The null is not sharp since the restriction set is a strict subset of the treatment assignment space.